OpenStudy (anonymous):

The graph of y = 2^x and the graph of y = 2^-x are symmetrical. What is the line of symmetry?

OpenStudy (anonymous):

The equation of the line of symmetry is x = -b/2a So one way to do this is put your equation into ax^2 + bx + c form: So expand: y = -2(x^2 + -2x + 1) +3 Now simplify: y = -2x^2 + 4x +1 Now you have a= -2, b=4, c=1 So plugging these into the equation for the line of symmetry: x = -4/2(-2) x = 1 is your equation. Ignore the previos answers.

OpenStudy (anonymous):

My options are as follows: the x-axis y = x the y-axis y = -x